OpenStudy (summersnow8):

1) Calculate the pH of a buffer solution prepared by mixing 60.0 mL of 1.00 M lactic acid and 25.0 mL of 1.00 M sodium lactate. (The pKa for lactic acid is 3.86. The molar mass for lactic acid is 90.1 g/mol and the molar mass for sodium lactate is 112.1 g/mol.) 2) a. If you mix equal volumes of 0.1 M HCl and 0.2 M PIPES (base form), is the resulting solution a good buffer? (The pKa of PIPES is 6.80.) b. Why or why not? -The amount of PIPES is much higher than the amount of HCl added. -PIPES (base form) only works as a buffer when mixed with an equal number of moles of its conjugate acid. -The concentrations of the free acid and base forms of PIPES are equal under these conditions. -An equal volume of 0.1 M NaOH rather than HCl should have been added to the PIPES base. 3) If you have 100. mL of a 0.150 M HEPES buffer at pH 7.55 and you add 2.00 mL of 1.00 M HCl, what will be the new pH? (The pKa of HEPES is 7.55.) @photon336 @whpalmer4

OpenStudy (kevin):

I'll saved this question for tomorrow if there are no one answered. I'm too sleepy to answer this. Lol

OpenStudy (cuanchi):

for calculate the pH of a buffer use the henderson hasselbalch equation. You have to calculate the final concentrations of the base (salt) and the acid in the mix before putting these values in the formula

OpenStudy (summersnow8):

This is my work:

OpenStudy (summersnow8):

This was my attempt at #3 @cuanchi

OpenStudy (cuanchi):

I see... you dont need to do the subtraction of the moles of salt and acid. Also because the acid and the salt are the same concentration you can use just the volumes in the relationship. At the end the relationship in volume is going to be equal to the relationship of moles or to the relationship of concentration acid= 0.06moles/ 0.085 L = 0.706M base= 0.025moles/0.085 L= 0.294 M 25/60= 0.025/0.060= 0.29/0.706 = 0.417 pH= 3.86 + log 0.417 pH= ????

OpenStudy (summersnow8):

pH = 3.48

OpenStudy (cuanchi):

I think so... are you sure?

OpenStudy (summersnow8):

That is what I plugged into the calculator

OpenStudy (cuanchi):

analyze your answer!!! are you expecting a pH higher or lower than the pKa? do you have more acid or more salt in the buffer?

OpenStudy (summersnow8):

I would expect the solution to be more acidic because more lactic acid was added. pKa indicates a weak acid....

OpenStudy (cuanchi):

yes makes sense

OpenStudy (kevin):

1) pH = pKa – log acid/ base pH = 3.86 – log (60/25) pH = 3.86 – 0.38 pH = 3.48 2) a. yes, because they have similar molarity b. - no, it’s higher but not much. - no. If it has equal number with the acid, it will no have mol left. Then it would be called hydrolisis solution. - I don’t know, but I guess it needs big coincidence to make it happen. So my answer is no. - No. If we want to make a buffer then will must added weak base + strong acid or vice versa. We cannot make buffer with strong base + weak base. 3) pH = pKa – log acid/ base pH = 7.55 – log (2/15) pH = 7.55 – 0.88 pH = 6.67 So, it’s decrease amount of 0.88